A solution contains 0.11 M sodium chloride and 0.11 M potassium fluoride. Solid lead nitrate is added slowly to this mixture. What ionic compound precipitates first from the solution? Solubility product constant data is found in the Chemistry References.
Answer #1

[Pb2+] = Ksp (PbCl2)/[Cl-]2 = 1.7×10-5/(0.11M)2

= 1.405×10-3 M = minimum concentration of Pb2+ require to precipitate PbCl2

[Pb2+] = Ksp (PbF2)/[F-]2 = 2.05×10-8/(0.11M)2

= 1.69×10-6 M = minimum concentration of Pb2+ required to precipitate PbF2 .

Answer: PbF2 (lead (II) fluoride ) will precipitate first .